OpenStudy (anonymous):

can someone teach me how to do these?? solve the nonlinear system of equations. {x^2+y^2=144 {x^2-4y^2=64

OpenStudy (anonymous):

How good are you at solving the linear versions of these?

OpenStudy (anonymous):

{x^2+y^2=144| {x^2-4y^2=64| *-1 ----------------- + 5y^2=80 y^2=16 y=+-4 now substitute each of this values into some equation and find the corresponding x values

OpenStudy (anonymous):

okay...thanks for your help myko!